# From Propeller Damage Estimation and Adaptation to Fault Tolerant Control: Enhancing Quadrotor Resilience

### v1

@@ Abstract
Aerial robots are required to remain operational even in the event of system disturbances, damages, or failures to ensure resilient and robust task completion and safety. One common failure case is propeller damage, which presents a significant challenge in both quantification and compensation. We propose a novel adaptive control scheme capable of detecting and compensating for multi-rotor propeller damages, ensuring safe and robust flight performances. Our control scheme includes an L1 adaptive controller for damage inference and compensation of single or dual propellers, with the capability to seamlessly transition to a fault-tolerant solution in case the damage becomes severe. We experimentally identify the conditions under which the L1 adaptive solution remains preferable over a fault-tolerant alternative. Experimental results validate the proposed approach, demonstrating its effectiveness in running the adaptive strategy in real time on a quadrotor even in case of damage to multiple propellers.
@@ I Introduction
Micro Aerial Vehicles (MAVs) such as quadrotors are becoming ubiquitous in applications such as search and rescue [1] and aerial photography [2]. However, MAV-related accidents hinder the growth of the industry and erode public confidence in drone safety. As a result, it is necessary to develop inference and control approaches that can guarantee safe and reliable flight in case of system damage. Propellers on MAVs are susceptible to damage, especially in the event of collisions or after prolonged use, primarily due to their size, location, and lightweight construction. In addition, detecting the occurrence and location of propeller failures is challenging and difficult to estimate. Indirectly sensing individual motor thrusts is further complicated due to the nonlinear dynamics of the quadrotor and the difficulty of measuring higher order angular acceleration or moment terms. Our work specifically targets quadrotors and introduces an adaptive control scheme to autonomously infer and adjust quadrotor flight behavior in response to propeller damages (defined in our case as the percent of thrust loss from a damaged propeller compared to a healthy propeller spinning at the same speed) or failures. This is achieved by first using adaptive control to adapt to damage and then inferring the damage from the adaptive control action unlike other works which first estimate a motor damage and then perform a corrective action.
@@
This work makes the following contributions. First, we introduce an adaptive control scheme in case of propeller damage, ensuring robust flight performances. Our solution combines an L1 adaptive control scheme, capable of detecting and compensating for single or dual propeller damages with minimal computational overhead and low latency, along with a fault-tolerant mode. This mode can be seamlessly activated when propeller damage reaches a critical level. Second, we study the trade-offs between adaptive control and fault-tolerant control as shown in Fig. 1 and experimentally identify the best operating modality according to the level of propeller damage. Third, we conduct a comprehensive analysis in both simulation and real-world experiments to assess the extent to which the L1 adaptive controller can improve tracking performance under various degrees of propeller damage. Finally, we validate our strategy on a real quadrotor, successfully adapting to propeller damage and transitioning in real-time to a fault-tolerant solution. This is the first work to tackle precise estimation of the propeller damage which has significant more difficulties than motor damage estimates of other works. These difficulties include unequal loss between thrust and torque coefficients along with additional noise on the system from damaged propellers.
@@ II Related Works
Research in fault estimation covers a variety of vehicles detailed in this survey [3] such as fixed wings and rotary aerial robots. We can divide methods handling propeller damage into two categories: empirical and state estimation methods. Empirical works address propeller damage estimation through parameter identification [4] and sensor noise analysis [5]. Work in [4] successfully estimates the loss of effectiveness of a single motor but is only tested in simulation. The correlation between changes in accelerometer noise to variations in propeller speeds is used in [5] to estimate propeller damage, but the technique is only tested on a minuscule amount of propeller damage.
@@
Other works rely on state estimation to accurately measure the damage [6, 7]. These typically have a slower response time than empirical methods, and are only able to handle motor damage estimation rather than propeller damage as in the proposed work. This means they assume the damage from the propeller motor pair can be represented as a scaled factor of each individual RPM. For the case of propeller damage, this assumption is not accurate. The thrust and torque coefficient scales with the prop radius, $r$, as a function of $r^{4}$ and $r^{5}$ respectively [8]. This increases the complexity of our propeller damage compared to [6, 7] problem. Typically a reduced state Kalman Filter [9, 6, 7] is used in either a single [9, 6] or cascaded[10, 11] form. These studies test their systems by artificially injecting a fault that decelerates the motor rather than testing true propeller damages. State estimation through cascaded Kalman Filters can infer and adapt to $60\%$ motor damage in [11]. However, it is only tested in simulation for motor not propeller damage and does not include the possibility to switch to the fault-tolerant mode. Our work negates up to $40-60\%$ propeller damage in real-world experiments as shown in Fig. 1.
@@
Once propeller damage is estimated, there are two options to incorporate damage compensation into control. The first option is adaptation where the damaged propeller is given a scaled up motor action [6, 7] to compensate for the damage. This adaptation is effective for small and medium damages but is unable to stabilize the robot for severe damages. The second option is to employ a fault-tolerant control strategy that completely disables the damaged propeller [4, 12], which is typically employed in case of severe damage. However, fault-tolerant control requires the robot to surrender yaw control causing rapid spins [13, 14], making it unsuitable for small or medium damages. Our proposed work combines both methods to estimate and adapt to small or medium propeller damage, and transition to fault-tolerant for severe damage.
@@
Two notable adaptive techniques are Incremental Nonlinear Dynamic Inversion (INDI) [14] and L1[15]. These works provide additional control actions that augment the capabilities of a baseline controller to reject damages or disturbances without calculating propeller damage estimates. However neither work provides an estimate to judge when propeller damage is too severe for adaption. Other methods exist for adaptation [16, 17, 18], through model identification or automatic gain tuning, but cannot adapt to rapid changes like the complimentary above. In this work, we design a L1 adaptive technique to compensate for propeller damage and give accurate propeller damage estimates. Unlike INDI[14] that requires angular acceleration which is difficult to estimate, our method only requires angular velocity which can be obtained from an on board Inertial Measurement Unit.
@@
Existing fault-tolerant control strategies are adept at ensuring robust control even in the event of single or multiple rotor failures [13, 14, 19, 3]. However, these solutions require identifying the faulty propeller, and sacrificing yaw control, resulting in a rapid yaw spinning as depicted in Fig. 1. Our work introduces an L1 adaptive control scheme designed to detect, infer, and compensate for propeller damage, with the capability to smoothly transition to a fault-tolerant solution when severe damage is detected.
@@ III-A Preliminaries
$\begin{split}\dot{\mathbf{p}}&=\mathbf{v},\\ m\dot{\mathbf{v}}&=mg\mathbf{e}_{3}+f\mathbf{R}\mathbf{e}_{3},\\ \dot{\mathbf{q}}&=\frac{1}{2}\mathbf{q}\otimes\begin{bmatrix}0\\ \mathbf{\Omega}\\ \end{bmatrix},\\ \mathbf{M}&=\mathbf{J}\dot{\mathbf{\Omega}}+\mathbf{\Omega}\times\mathbf{J}\mathbf{\Omega},\end{split}$ (1)
where $\mathbf{p}\in\mathbb{R}^{3}$, and $\mathbf{v}\in\mathbb{R}^{3}$ are the position and velocity in the body frame. $\mathbf{R}\in SO(3)$ is a rotation matrix from the body fixed frame to the inertial frame, $\mathbf{q}$ is the quaternion representation of $\mathbf{R}$, and $\otimes$ is quaternion multiplication. $f$, $m$, $g$ are the thrust, mass and gravity respectively, $\mathbf{\Omega}\in\mathbb{R}^{3}$ is the angular velocity with respect to the body frame, $\mathbf{M}=\begin{bmatrix}M_{1}&M_{2}&M_{3}\end{bmatrix}^{\top}$ are the moments around the three body frame axes, and $\mathbf{J}\in\mathbb{R}^{3\times 3}$ is the robot inertia matrix.
Figure 3: Our Resilient Control Scheme.
@@ III-B Control Design
We adapt the geometric controller derived in [20] for the possibility of propeller damage, or losing one or two opposing rotors in flight. An outer controller solves for the desired attitude $\mathbf{R}$, thrust $f$, and angular velocities $\mathbf{\Omega}$. The inner controller solves for the moments required to control the attitude of the MAV. The geometric controller tracks position $\mathbf{p}$ and yaw $\psi$ by generating force and moments control signals. The L1 augmentation is inserted into the controller as seen in Fig. 3. The L1 controller uses a state predictor to estimate the partial state $[\mathbf{v}^{\top}\leavevmode\nobreak\ \mathbf{\Omega}^{\top}]^{\top}$ with the unknown disturbances. The adaptation law compares the prediction to the measured state and generates the control output to compensate for disturbances. If the adaptation law accurately estimates the unknown disturbances then the tracking error will converge to zero [21] assuming the augmented action is with in the actuator limits.
@@ III-B1 Outer-loop Controller
As per the standard geometric control algorithm, the thrust vector is normalized and chosen as the third body axis, or thrust vector $\mathbf{b}_{3}$
@@
Next, we formulate the rotation matrix based on $\psi$ or the yaw of the vehicle. We denote the individual elements of the thrust vector as $\mathbf{b}_{3}=\begin{bmatrix}b_{3x}&b_{3y}&b_{3z}\end{bmatrix}^{\top}$. We can formulate a quaternion representing the tilt, $\mathbf{q}_{tilt}$, and yaw rotation $\mathbf{q}_{\psi}$ such as in [22]. The below quaternion formulations are all scalar first notation.
@@
Multiplying the quaternion $\mathbf{q}_{tilt}$ and $\mathbf{q}_{\psi}$ we formulate the full vehicle’s rotation with $\mathbf{q}$ being the quaternion representation of $\mathbf{R}$. A rotation matrices representation of eq. (3) is in [13]
$\mathbf{q}=\mathbf{q}_{tilt}\otimes\mathbf{q}_{\psi}.$ (4)
@@
where $\mathbf{p}_{d}$ and $\mathbf{v}_{d}$ are desired position and velocity of the vehicle and $\mathbf{k}_{p},\mathbf{k}_{v}\in\mathbb{R}^{3}$ are the gains for the respective errors. Lastly, the angular velocity $\mathbf{\Omega}$, is solved using
$\begin{bmatrix}0\\ \mathbf{\Omega}\\ \end{bmatrix}=2\mathbf{q}^{-1}\otimes\dot{\mathbf{q}},$ (6)
as a function $\dot{\mathbf{b}}_{3}$ and $\dot{\psi}$ such as in [22] .
@@ III-B2 Inner-loop Controller
$\mathbf{e}_{R}=\frac{1}{2}(\mathbf{R}_{d}^{\top}\mathbf{R}-\mathbf{R}^{\top}\mathbf{R}_{d})^{\vee}.$ (7)
@@
where $\mathbf{b}_{3d}$ is the normalized desired thrust vector and $\mathbf{b}_{3}$ is the $z$ axis of the body fixed frame we get from the localization. The cross product of $\mathbf{b}_{3d}$ and $\mathbf{b}_{3}$ measures how much the actual direction deviates from the desired direction. The angular velocity error, $\mathbf{e}_{\Omega}$ is calculated by
@@ III-C Adaptive Control
We supplement our controller with an L1 adaptive controller by including two elements: 1)1)1 ) a state predictor to estimate the evolution of the partial state and the disturbances present and 2)2)2 ) an adaption law to solve for an action to counter the error between the estimated and true partial state.
III-C1 State Predictor
We adopt a partial state predictor to estimate how the partial state $\mathbf{z}=\begin{bmatrix}\mathbf{v}^{\top}&\mathbf{\Omega}^{\top}\end{bmatrix}^{\top}$ evolves given an input $\bm{\mu}=\begin{bmatrix}f&\mathbf{M}^{\top}\end{bmatrix}^{\top}$ and disturbances $\bm{\sigma}\in\mathbb{R}^{6}$. These disturbances $\sigma=\begin{bmatrix}\mathbf{f}_{\text{ext}}^{\top}&\mathbf{M}_{\text{ext}}^{\top}\end{bmatrix}^{\top}$ represent the additional force and moments on the vehicle both expressed in the body frame not accounted in the standard controller dynamic model. Let $k$ denote the discrete time step of the system. To acquire the next state from derivatives we simply perform an Euler approximation given a time step $dt$

### v2

@@ Abstract
Aerial robots are required to remain operational even in the event of system disturbances, damages, or failures to ensure resilient and robust task completion and safety. One common failure case is propeller damage, which presents a significant challenge in both quantification and compensation. In this paper, we propose a novel adaptive control scheme capable of detecting and compensating for multi-rotor propeller damages, ensuring safe and robust flight performances. Our solution combines an L1 adaptive controller with an optimization routine for damage inference and compensation of single or dual propellers, with the capability to seamlessly transition to a fault-tolerant solution in case the damage becomes severe. We experimentally identify the conditions under which the L1 adaptive solution remains preferable over a fault-tolerant alternative. Experimental results validate the proposed approach demonstrating the ability of our solution to adapt and compensate onboard in real time on a quadrotor for damages even when multiple propellers are damaged.
Index Terms:
Aerial Systems: Mechanics and Control; Aerial Systems: Applications
@@ I Introduction
Micro Aerial Vehicles (MAVs) such as quadrotors are becoming ubiquitous in applications such as search and rescue and aerial photography [1]. However, MAV-related accidents hinder the growth of the industry and erode public confidence in drone safety. As a result, it is necessary to develop inference and control approaches that can guarantee safe and reliable flight in case of system damage. Propellers on MAVs are susceptible to damage, especially in the event of collisions or after prolonged use, primarily due to their size, location, and lightweight construction. In addition, detecting the occurrence and location of propeller failures is challenging. Indirectly sensing individual motor thrusts is further complicated due to the nonlinear dynamics of the quadrotor and the difficulty of measuring higher order angular acceleration or moment terms. Our work specifically introduces an adaptive control scheme for quadrotors to autonomously infer and adjust flight behavior in response to propeller damages or failures.
@@
This work makes the following contributions. First, we propose a propeller damage estimation technique by combining an L1 Adaptive algorithm and an optimization routine with no need for direct RPM (Rotations Per Minute) measurements or estimation of angular accelerations, making it suitable for low-cost systems. Second, we propose a holistic control scheme coupling the L1 adaptive and fault-tolerant controller modules for autonomous transition from L1 adaptation to the fault-tolerant mode in situations where adaptation alone proves to be insufficient for ensuring safe and effective flight. This approach enhances the system’s capability to manage the full range of propeller damage. Finally, we demonstrate the performance of the damage estimation and controller design in extensive real-world experiments and compare our estimation method to using a standard PID (Proportional Integral Derivative) controller. This is the first work to tackle precise estimation and compensation of propeller damage rather than considering a reduction of effectiveness of a motor-propeller pair. Challenges include unequal loss between thrust and torque coefficients along with additional noise on the system from damaged propellers as well as the unavailability of direct RPM measurements.
@@ II Related Works
Damage Estimation. Research in fault estimation covers a variety of vehicles detailed in this survey [2] such as fixed wings and rotary aerial robots. We can divide methods handling propeller damage into two categories: empirical and state estimation methods. Empirical works address propeller damage estimation through parameter identification [3] or sensor noise analysis [4]. The approach proposed in [3] successfully estimates the loss of effectiveness of a single motor but is only tested in simulation, whereas [4] uses accelerometer noise to estimate propeller damage, but is only tested on a minuscule amount of propeller damage.
@@
Other works rely on state estimation to accurately measure the damage [5, 6]. These typically have a slower response time than empirical methods, and are only able to handle motor damage estimation rather than propeller damage as in the proposed work. They typically assume the damage from the propeller motor pair can be represented as a scaled factor of each individual RPM. For the case of propeller damage, this assumption is not accurate. The thrust and torque coefficients scale with the propeller radius, $r$, as a function of $r^{4}$ and $r^{5}$ respectively [7]. This increases the complexity of our propeller damage compared to the problems in [5, 6, 8]. Typically a reduced state Kalman Filter [9, 5, 6] is used in either a single [9, 5] or cascaded[10, 11] form. These studies test their systems by artificially injecting a fault that decelerates the motor rather than testing true propeller damages. State estimation through cascaded Kalman Filters can infer and adapt up to $60\%$ motor damage in [11]. However, it is only tested in simulation for motor damage and does not include the possibility to switch to the fault-tolerant mode. Our work negates up to $40-60\%$ propeller damage in real-world experiments as shown in Fig. 1. Uniquely, [8] implements a Kalman filter method to detect a complete loss of a propeller and is able to detect a fault within $0.1\leavevmode\nobreak\ $\mathrm{s}$$ of occurrence. However, this method requires accurate RPM measurements and cannot adapt to propeller damage.
@@
Damage Compensation. Once propeller damage is estimated, there are two options to incorporate damage compensation into control. The first option is adaptation where the damaged propeller is given a scaled up motor action [5, 6] to compensate for the damage. This adaptation is effective for small and medium damages but is unable to stabilize the robot for severe damages. The second option is to employ a fault-tolerant control strategy that completely disables the damaged propeller [3, 12] but sacrifices yaw control.
@@
A few adaptive techniques are Incremental Nonlinear Dynamic Inversion (INDI) [13], dual loop disturbance observers [14], and L1 [15, 16, 17]. These works provide additional control actions that augment the capabilities of a baseline controller to reject some actuator damages or disturbances. However, none of the above works provides an estimate to judge when propeller damage is too severe for adaption as shown in the proposed work. Other methods exist for adaptation [18, 19, 20], through model identification or automatic gain tuning, but cannot adapt to rapid changes. In this work, we design an L1 adaptive technique to compensate and accurately estimate propeller damage. Unlike INDI[13] that requires angular acceleration which is difficult to estimate, our method only requires angular velocity which can be obtained from an on board Inertial Measurement Unit (IMU).
@@
Existing fault-tolerant control strategies are adept at ensuring robust control even in the event of single or multiple rotor failures [21, 13, 2, 22]. However, these solutions require identifying the faulty propeller, and sacrificing yaw control which results in a rapid yaw spin as depicted in Fig. 1 making them unsuitable for small to medium damages. Our work introduces a holistic control scheme designed to detect, infer, and compensate for propeller damage, with the capability to smoothly transition to a fault-tolerant solution when severe damage is detected allowing effective operations for safe recovery and navigation.
@@ III-A Preliminaries
$\begin{split}\dot{\mathbf{p}}&=\mathbf{v},\\ m\dot{\mathbf{v}}&=mg\mathbf{e}_{3}+\mathbf{q}\otimes f\mathbf{e}_{3},\\ \dot{\mathbf{q}}&=\frac{1}{2}\mathbf{q}\otimes\begin{bmatrix}0\\ \mathbf{\Omega}\\ \end{bmatrix},\\ \mathbf{M}&=\mathbf{J}\dot{\mathbf{\Omega}}+\mathbf{\Omega}\times\mathbf{J}\mathbf{\Omega},\end{split}$ (1)
where $\mathbf{p}\in\mathbb{R}^{3}$, and $\mathbf{v}\in\mathbb{R}^{3}$ are the position and velocity in the world frame. $\mathbf{R}\in SO(3)$ is a rotation matrix from the body fixed frame to the inertial frame, $\mathbf{q}$ is the quaternion representation of $\mathbf{R}$, and $\otimes$ is quaternion multiplication. $f$, $m$, $g$ are the thrust, mass and gravity respectively, $\mathbf{\Omega}\in\mathbb{R}^{3}$ is the angular velocity with respect to the body frame, $\mathbf{M}=\begin{bmatrix}M_{1}&M_{2}&M_{3}\end{bmatrix}^{\top}$ are the moments around the three body frame axes, and $\mathbf{J}\in\mathbb{R}^{3\times 3}$ is the robot inertia matrix.
Figure 3: The cascaded control scheme, shows the L1 Adaptation Controller (blue), receives the partial state feedback from the system. The Fault Tolerant Controller (pink) is activated only once damage estimate exceeds $50\%$.
@@ III-B Control Design
We adapt the geometric controller derived in [23] for the possibility of propeller damage, or fully losing one or two opposing rotors in flight. An outer controller solves for the desired attitude $\mathbf{R}$, thrust $f$, and angular velocities $\mathbf{\Omega}$. The inner controller solves for the moments required to control the attitude of the MAV. The geometric controller tracks position $\mathbf{p}$ and yaw $\psi$ by generating force and moments control signals. The L1 augmentation is inserted into the controller as seen in Fig. 3. The L1 controller uses a state predictor to estimate the partial state $[\mathbf{v}^{\top}\leavevmode\nobreak\ \mathbf{\Omega}^{\top}]^{\top}$ with the unknown disturbances. The adaptation law compares the prediction to the measured state and generates the control output to compensate for disturbances. If the adaptation law accurately estimates the unknown disturbances then the tracking error will converge to zero [24] assuming the augmented action is with in the actuator limits.
@@ III-B1 Outer-loop Controller
As per the standard geometric control algorithm, the thrust vector is normalized and chosen as the third body axis, or thrust vector $\mathbf{b}_{3}$ as
@@
Next, we formulate the rotation matrix based on $\psi$ or the yaw of the vehicle. We denote the individual elements of the thrust vector as $\mathbf{b}_{3}=\begin{bmatrix}b_{3x}&b_{3y}&b_{3z}\end{bmatrix}^{\top}$. We can formulate a quaternion representing the desired tilt, $\mathbf{q}_{tilt}$, and yaw rotation $\mathbf{q}_{\psi}$ such as in [25]. The below quaternion formulations are all scalar first notation.
@@
Multiplying the quaternion $\mathbf{q}_{tilt}$ and $\mathbf{q}_{\psi}$ we formulate the full vehicle’s rotation with $\mathbf{q}_{d}$ being the quaternion representation of desired orientation, $\mathbf{R}_{d}$. A rotation matrices representation of eq. (3) is in [21]
$\mathbf{q}_{d}=\mathbf{q}_{tilt}\otimes\mathbf{q}_{\psi}.$ (4)
@@
where $\mathbf{p}_{d}$ and $\mathbf{v}_{d}$ are desired position and velocity of the vehicle and $\mathbf{k}_{p},\mathbf{k}_{v}\in\mathbb{R}^{3}$ are the gains for the respective errors. Lastly, the desired angular velocity $\mathbf{\Omega}_{d}$, is
$\begin{bmatrix}0\\ \mathbf{\Omega}_{d}\\ \end{bmatrix}=2\mathbf{q}_{d}^{-1}\otimes\dot{\mathbf{q}_{d}},$ (6)
as a function of the desired $\dot{\mathbf{b}}_{3}$ and $\dot{\psi}$ such as in [25] .
@@ III-B2 Inner-loop Controller
$\mathbf{e}_{R}=\frac{1}{2}(\mathbf{R}_{d}^{\top}\mathbf{R}-\mathbf{R}^{\top}\mathbf{R}_{d})^{\vee},$ (7)
@@
where $\mathbf{b}_{3d}$ is the normalized desired thrust vector and $\mathbf{b}_{3}$ is the current $z$ axis of the body fixed frame we get from the localization. The cross product of $\mathbf{b}_{3d}$ and $\mathbf{b}_{3}$ measures how much the actual direction deviates from the desired direction. The angular velocity error, $\mathbf{e}_{\Omega}$ is calculated by
@@ III-C Adaptive Control
We supplement our controller with an L1 adaptive controller by including two elements: 1)1)1 ) an adaption law to solve for an action to counteract the un-modeled disturbances on the system, and 2)2)2 ) a state predictor to estimate a predicted state, linear $\mathbf{v}_{p}$ and angular $\mathbf{\Omega}_{p}$ velocities given the L1 action and un-modeled disturbances.
III-C1 L1 Adaptation Law
Algorithm 1 L1 Adaptation Control Step $[k]\rightarrow[k+1]$
0:   Measured velocities $\mathbf{v}_{m}[k],\mathbf{\Omega}_{m}[k]$
0:   L1 Force and Moment $f_{L1}[k],\mathbf{M}_{L1}[k]$
0:   Predicted velocities $\mathbf{v}_{p}[k],\mathbf{\Omega}_{p}[k]$
0:
1:  Calculate the Disturbances, $\bm{\sigma}[k]$ (eq. (11))
2:  $f_{L1}[k]=\text{LPF}(-\bm{\sigma}_{3}[k])$
3:  $\mathbf{M}_{L1}[k]=\text{LPF}(-\bm{\sigma}_{4:6}[k])$
3:
4:  Solve for $\dot{\mathbf{v}}_{p}[k]$ (eq. (12)) and $\dot{\mathbf{\Omega}}_{p}[k]$ (eq. (13))
5:  Solve for $\mathbf{v}_{p}[k+1]$ and $\mathbf{\Omega}_{p}[k+1]$ (eq. (14))
6:  Update the Current $[k]$ with the Next State $[k+1]$
7:  Add $f_{L1}[k]$ the base thrust control eq. (5)
8:  Add $\mathbf{M}_{L1}[k]$ to the base moment control eq. (10)
First, we characterize the disturbances $\bm{\sigma}=\begin{bmatrix}\mathbf{f}_{\text{ext}}^{\top}&\mathbf{M}_{\text{ext}}^{\top}\end{bmatrix}^{\top}$ which represents the un-modeled force and moments on the vehicle both expressed in the body frame. These disturbances are obtained using the error between the predicted state velocity $\begin{bmatrix}\mathbf{v}_{p}^{\top}&\mathbf{\Omega}_{p}^{\top}\end{bmatrix}$ generated from the state predictor in Section III.C.2 and measured velocities, $(\mathbf{v}_{m},\mathbf{\Omega}_{m})$. Let $k$ denote the discrete time step of the system. We set a diagonal gain defined by the user $\mathbf{K}\in\mathbb{R}^{6\times 6}$ to generate a specific gain, $\mathbf{A}$, based on the time step, $dt$, between $k$ and $k+1$ time instants with